# An Efficient Finite Difference Approximation via a Double Sample-Recycling Approach

### v1

@@ 4.1 Constructing the DSR-CFD estimator
  • Variance reduction: For any $k=1,...,K$, $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is highly positively correlated with $\Delta_{j}(h_{k})$ since $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is an approximation of $\Delta_{j}(h_{k})$. Although $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is not zero-mean, it can be seen as a control variate because only the variance is considered. In the end, these control variates will reduce the variance (see Glasserman, 2013).
  • Possible bias reduction: As observed in (15), the regression problem is non-canonical and $\mathbb{E}[\mathcal{E}_{e}]\neq 0$. Therefore, $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ is a biased estimator of $\mathbb{E}[\Delta(h_{k})]$ which leads to the difference between the biases. It is more obvious if we rewrite the DSR-CFD estimator as
    $\widehat{\Delta}_{n,\widehat{h}_{n}}=$ $\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+n_{b}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\widehat{\Delta}_{n_{b},h_{k}}-\mathbb{E}\Delta(h_{k})\right)+Kn_{b}\left(1,{\widehat{h}_{n}}^{2}\right)\widehat{\bm{\beta}}_{e}\right)$
    $+\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\mathbb{E}\Delta(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)$
    $=$ $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}+\frac{n_{b}}{n}\sum_{k=1}^{K}\frac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\mathbb{E}\Delta(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right).$
    Note that this bias adjustment depends on $D$ (see (15)) while the bias of the optimal CFD estimator depends on $B$ (see (9)). If these two bias terms have a tendency to cancel each other out, the increment of $\mathrm{Bias}\left[\widehat{\Delta}_{n,\widehat{h}_{n}}\right]$ may pull $\mathrm{Bias}\left[\widehat{\Delta}_{n,h^{*}}\right]$ towards 0, thus reducing the bias.
@@
In Section 6, we will present several numerical examples to validate the aforementioned properties of our proposed DSR-CFD estimator.
@@ 5 Implementation
Previous sections have provided efficient estimators for $\bm{\beta}_{e}$ and $\bm{\beta}_{v}$ and constructed a double sample-recycling based estimator. The theoretical bias, variance and MSE have also been calculated, that is Theorem 4. This section will provide the details for estimating $\bm{\beta}_{e}$ and $\bm{\beta}_{v}$ and constructing the DSR-CFD estimator.
5.1 Estimating parameters $B$ and $\sigma^{2}(\theta_{0})$
In this subsection, we present the specific steps for estimating unknown constants using the bootstrap method as described in Section 3. We generate i.i.d. $c_{1},...,c_{K}$ from $\mathcal{P}_{0}$ and for $k=1,...,K$, let $h_{k}=c_{k}n_{b}^{-1/10}$. At each perturbation $h_{k}$, we generate $n_{b}$ samples $\bm{\Delta}(h_{k})=\{\Delta_{1}(h_{k}),...,\Delta_{n_{b}}(h_{k})\}$. Repeat independently the bootstrap procedure in Section 3 for $I$ times, we obtain $I$ versions of $\widehat{\Delta}^{b}_{n_{b},h_{k}}$, denoted by $\left\{\widehat{\Delta}^{b}_{n_{b},h_{k}}(1),...,\widehat{\Delta}^{b}_{n_{b},h_{k}}(I)\right\}$. Then $\mathbb{E}_{*}\widehat{\Delta}_{n_{b},h_{k}}$ is approximated by
@@ 5.1 Estimating parameters $B$ and $\sigma^{2}(\theta_{0})$
It is worth noting that both of the regression problems (14) and (16) exhibit heteroscedasticity. This arises from the fact that elements of $\mathcal{E}_{e}$ in (15) and $\mathcal{E}_{v}$ in (17) possess varying variances, due to different values of $h_{k}$’s. The primary approach to addressing heteroscedasticity is through the weighted least squares method. This involves determining appropriate weights to eliminate or mitigate the heteroscedasticity in the regression models, i.e., (14) and (16). As the estimator obtained through the weighted least squares method exhibits the same convergence rate as that derived from the ordinary least squares method (Hayashi, 2011), we omit an in-depth analysis of the weighted least squares method.
@@
However, when estimating $\sigma^{2}(\theta_{0})$, we lack such standard deviations. From $\mathcal{E}_{e}$ in (17), it can be seen that the heteroscedasticity of the regression model (16) arises from the different variances of $\phi(\cdot)$ for different values of $h_{k}$’s. It follows from (13) that $h_{k}^{2}\phi(h_{k})$ ($k=1,...,K$) share the same variance. Therefore, we can handle the heteroscedasticity resulting from differences in perturbations by multiplying both sides of (16) by $h_{k}^{2}$ ($k=1,...,K$) simultaneously. That is, we regress $\widetilde{\bm{Y}}_{v}$ on $\widetilde{\bm{X}}_{v}$, where
@@ 5.2 Estimating $\alpha^{\prime}(\theta_{0})$
Once we have generated $\Delta_{j}(h_{k})$, where $j=1,2,...,n_{b}$ and $k=1,2,...,K$, and have estimated the unknown constants using these samples, we can transform these samples using (31) and construct the DSR-CFD estimator using (32). We summarize the above procedure as Algorithm 1:
@@
    Usually, we choose $\mathcal{P}_{0}$ as a truncated normal distribution $\mathcal{P}_{0}=\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$, where $0<L<U<\infty$. This setting ensures that $\mathbb{E}[H_{K}]$, $\mathrm{Var}[H_{K}]$, $\mathbb{E}[V_{K}]$, $\mathbb{E}\left[\widetilde{H}_{K}\right]$, $\mathrm{Var}\left[\widetilde{H}_{K}\right]$, $\mathbb{E}\left[\widetilde{V}_{K}\right]$, $\mathbb{E}\left[\widehat{H}_{K}\right]$,$\mathrm{Var}\left[\widehat{H}_{K}\right]$ and $\mathbb{E}\left[\widehat{V}_{K}\right]$ are all bounded. Note that for $c_{1},...,c_{K}\sim\mathcal{P}_{0}$, the condition $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ holds only when $U\leq\sqrt{2}L$, which is a stringent condition. However, we have mentioned that $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ is a sufficient condition of $\left|\left|{\rm{Diag}}(\bm{c^{-1}})\bm{Pc}\right|\right|_{2}^{2}\leq K$ which is satisfied more easily for $\mathcal{P}_{0}$ with different parameters.
@@
    In step 1, the pilot samples are generated at the perturbations $h_{k}=c_{k}n_{b}^{-1/10}$, rather than $c_{k}n_{b}^{-1/6}$. This is because (23) and (24) indicate that in the pilot stage, the optimal perturbation is of order $n_{b}^{-1/10}$.
@@
    In step 2, we generate the remaining samples at $\widehat{h}_{n}$ defined as in (28), instead of $\widehat{h}_{n_{2}}$ defined as in (25), because we will utilize the total $n$ samples to construct the CFD estimator.
Input: The number of total sample pairs $n$, the number of perturbation parameters $K$, the number of pilot sample pairs at each perturbation $n_{b}$, the number of bootstrap $I$. Initial perturbation generator $\mathcal{P}_{0}$.
@@
    Regress $\widetilde{\bm{Y}}_{e}$ on $\widetilde{\bm{X}}_{e}$, where $\widetilde{\bm{Y}}_{e}$ and $\widetilde{\bm{X}}_{e}$ are defined as in (41), and the slope is $\widehat{B}$.
    Regress $\widetilde{\bm{Y}}_{v}$ on $\widetilde{\bm{X}}_{v}$, where $\widetilde{\bm{Y}}_{v}$ and $\widetilde{\bm{X}}_{v}$ are defined as in (42), and the slope is $\widehat{\bm{\beta}}_{v}$.
@@
Step 2. Construct the DSR-CFD estimator of $\alpha^{\prime}(\theta_{0})$.
@@
    Reuse $\Delta_{j}(h_{k})$, where $j=1,2,...,n_{b}$ and $k=1,2,...,K$. That is, calculate $\Delta_{j}^{t}(h_{k})=\dfrac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)+\left(1,{\widehat{h}}_{n}^{2}\right)\widehat{\bm{\beta}}_{e}$ for $j=1,2,...,n_{b}$, and $k=1,2,...,K$.
@@
    Denote the number of remaining sample pairs as $n_{2}=n-Kn_{b}$. Generate $n_{2}$ sample pairs at $\widehat{h}_{n}$ and calculate $\left\{\Delta_{1}(\widehat{h}_{n}),...,\Delta_{n_{2}}(\widehat{h}_{n})\right\}$.
@@
    The DSR-CFD estimator
    $\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+\sum_{k=1}^{K}\sum_{j=1}^{n_{b}}\Delta_{j}^{t}(h_{k})\right).$
@@ 6 Numerical Experiments
In this section, we use three examples to check the performance of our method. Examples 1 and 2 are from Li and Lam (2020). Example 3 is the generic $M/M/1$ queueing system from Lam et al. (2022). All numerical results are based on 1000 replications.
We will compare the performance of our proposed method and the EM-CFD method proposed by Li and Lam (2020). The detailed settings of our proposed method are presented as follows:
  • In the first and third examples, we let the initial perturbation generator $\mathcal{P}_{0}=\psi\left(0,1,0.1,\infty\right)$ ($\psi\left(0,0.1,0.01,\infty\right)$ in Example 2), the number of bootstrap $I=1000$. During the first phase, we set the number of regression points as $K=10$ or $20$.
  • The bootstrap-based method, labeled by “BOOT-CFD”, represents using the bootstrap method to estimate the unknown constants but not reusing the pilot samples.
@@
Consider estimating the first order derivative of the function $\alpha(\theta)=k{\rm sin}(\theta)$ at the point $\theta_{0}=0$, where $k$ is a priori unknown parameter.
@@
$\alpha^{\prime}(\theta_{0})=k,\quad B=-k/6,\quad\sigma^{2}(\theta_{0})=1,$
@@
Initially, we assess the validity and robustness of the bootstrap method for estimating unknown constants and perturbations. In the implementation, we set $k=10$, $K=10$ and utilize 200 pilot sample pairs. The results, presented in Figures 1 and 2, display three subfigures from left to right, representing the estimators of $B$, $\sigma^{2}(\theta_{0})$, and $h$ across 1000 replications. When executing the EM-CFD algorithm, to align with the setting in Li and Lam (2020), we generate $h_{1},\ldots,h_{200}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times 200^{-1/5}\right)$.
In the left subfigure of Figure 1, we observe that our method’s robustness in capturing third-order derivative information is slightly lower than that of the EM-CFD algorithm in case 1 since DSR-CFD method has a wider box. This happens because the DSR-CFD method has only $K=10$ regression points instead of 200. Nonetheless, the DSR-CFD method is still comparable due to the reduced error at each regression point. The middle subfigure of Figure 1 indicates our approach effectively estimates variance information, a challenging task for the EM-CFD algorithm as it incorporates the variance of the perturbation generator. The right subfigure provides a comprehensive view of perturbation estimation. Across all 1000 experiments, our method consistently outperforms the EM-CFD algorithm, approaching the true value more closely.
@@
Figure 1: Comparison of the estimates of the unknown constants and perturbation for case 1 in Example 1 $(k=10)$.
@@
In Figure 2, we illustrate the impact of various approaches on the estimation of $B$, $\sigma^{2}(\theta_{0})$, and $h$ in case 2 of Example 1, which deals with scenarios of heteroskedasticity. Analyzing the pair of subfigures presented on the left, it is apparent that our method captures the variance information effectively and provides a more robust estimate of $B$. From the right subfigure, we observe that our method yields a more accurate and reliable estimator of the perturbation. In contrast, the perturbations obtained by the EM-CFD algorithm deviate significantly from the true value, exhibit higher fluctuations.
@@
Figure 2: Comparison of the estimates of the unknown constants and perturbation for case 2 in Example 1 $(k=10)$.
In Figure 3, we depict the sensitivity of different methods to $r$, where the suffixes “1” and “2” denote cases 1 and 2, respectively, in Example 1. We set the number of sample pairs to 1000. When examining the EM-CFD method, we generate $h_{1},\ldots,h_{1000r}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times(1000r)^{-1/5}\right)$. As observed in the figure, the presence of heteroskedasticity complicates the problem for each method, resulting in inferior estimations. It is worth mentioning that the impact of heteroskedasticity is more pronounced in EM-CFD. When $r<0.55$, the EM-CFD method exhibits significant instability. The DSR-CFD method consistently outperforms other methods in various scenarios, surpassing the EM-CFD method in both simple and complex situations. Additionally, the DSR-CFD method demonstrates stability concerning $r$, while the performance of EM-CFD and BOOT-CFD worsens with increasing $r$. Even when $r=1$, the DSR-CFD method maintains strong performance, making it suitable for small sample sizes, where all samples are utilized for both constant and gradient estimation, effectively leveraging the sample information.
Figure 3: Illustration of the MSE of different methods with $r$ for cases 1 and 2 in Example 1 $(k=10)$.
@@
We consider estimating the first order derivatives of the polynomial function $\alpha(\theta)=1-6\theta+36\theta^{2}-53\theta^{3}+22\theta^{5}$ at different $\theta_{0}$. Assume that the observed variables at $\theta$ obey a normal distribution with mean $\alpha(\theta)$ and variance $0.05$, i.e., $Y(\theta)\sim\mathcal{N}(\alpha(\theta),0.05)$. From a simple calculation we obtain
$\alpha^{\prime}(\theta_{0})=-6+72\theta_{0}-159\theta_{0}^{2}+110\theta_{0}^{4},\quad B=-53+220\theta_{0}^{2},\quad\sigma^{2}(\theta_{0})=0.05,$
@@
Table 3 presents a comparison of bias, variance, and MSE between the DSR-CFD estimator (denoted by Bias-DSR, Var-DSR and MSE-DSR, respectively) and the Opt-CFD estimator (denoted by Bias-Opt, Var-Opt and MSE-Opt, respectively) across various sample sizes and estimation points (corresponding to different problems) in example 2. In Table 3, $B$ and $D$ are respectively the true values and Opt-CFD is meant to set $B$ and $\sigma^{2}(\theta_{0})$ as their true values, i.e., $B=-53+220\theta_{0}^{2}$ and $\sigma^{2}(\theta_{0})=0.05$. We take 50% of the total sample to estimate the unknown constants and set $K=10$ for DSR-CFD method.
Table 3: Comparison of the bias, variance and MSE of the DSR-CFD and Opt-CFD estimators in Example 2.
Budget $x$ $B$ $D$ Bias-DSR Bias-Opt Var-DSR Var-Opt MSE-DSR MSE-Opt
$10^{3}$ 0 $-53$ 22 $-0.1335$ $-0.0829$ 0.0117 0.0154 0.0295 0.0222
0.2 $-44.2$ 22 $-0.1356$ $-0.0789$ 0.0115 0.0149 0.0299 0.0211
0.4 $-17.5$ 22 $-0.1264$ $-0.0584$ 0.0105 0.0067 0.0265 0.0101
0.6 26.2 22 0.0170 0.0693 0.0102 0.0101 0.0105 0.0149
0.8 87.8 22 0.0658 0.1016 0.0163 0.0212 0.0206 0.0315
1 167 22 0.0991 0.1294 0.0217 0.0323 0.0315 0.0490
$10^{4}$ 0 $-53$ 22 $-0.0583$ $-0.0415$ 0.0021 0.0031 0.0055 0.0049
0.2 $-44.2$ 22 $-0.0558$ $-0.0339$ 0.0021 0.0031 0.0052 0.0042
0.4 $-17.5$ 22 $-0.0544$ $-0.0279$ 0.0017 0.0015 0.0047 0.0023
0.6 26.2 22 0.0099 0.0343 0.0018 0.0021 0.0019 0.0033
0.8 87.8 22 0.0340 0.0484 0.0033 0.0044 0.0044 0.0068
1 167 22 0.0549 0.0596 0.0055 0.0069 0.0085 0.0104
$10^{5}$ 0 $-53$ 22 $-0.0245$ $-0.0196$ 4.5424e-04 6.7200e-04 0.0011 0.0011
0.2 $-44.2$ 22 $-0.0222$ $-0.0179$ 4.2904e-04 6.3269e-04 9.2243e-04 9.5076e-04
0.4 $-17.5$ 22 $-0.0208$ $-0.0126$ 2.8572e-04 3.3534e-04 7.1882e-04 4.9412e-04
0.6 26.2 22 0.0088 0.0150 3.4222e-04 4.2576e-04 4.1963e-04 6.4935e-04
0.8 87.8 22 0.0200 0.0220 7.4610e-04 9.9544e-04 0.0011 0.0015
1 167 22 0.0270 0.0277 9.8655e-04 0.0015 0.0017 0.0023

### v4

@@ 4.2 Asymptotic Analyses of the Cor-CFD Estimator
  • Variance reduction: The second term of (38) is highly positively correlated with the second term of (37) because $\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}$ approximates $\widehat{\Delta}_{n_{b},h_{k}}$. Although the former is not mean 0, it can be viewed as a control variate of the latter. Therefore, the second term of (38) reduces the variance of $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$ to that of $\widehat{\Delta}_{n,\widehat{h}_{n}}$.
  • Possible bias reduction: The expectation of the second term of (38) depends on $D$ (see (15)) while the bias of $\widehat{\Delta}_{n,\widehat{h}_{n}}^{new}$ depends on $B$ (see (9)). If these two terms tend to cancel each other out, the overall bias will be reduced.
@@
In Section 6, we will present several numerical examples to validate the aforementioned properties of our proposed Cor-CFD estimator.
@@ 5 Implementation
This section provides the implementations for estimating $\bm{\beta}_{e}$ and $\bm{\beta}_{v}$ and constructing the Cor-CFD estimator.
5.1 Estimating $B$ and $\sigma^{2}(\theta_{0})$
In this subsection, we present the specific steps for estimating unknown constants using the bootstrap method as described in Section 3. We generate i.i.d. $c_{1},...,c_{K}$ from $\mathcal{P}_{0}$ and for $k=1,...,K$, let $h_{k}=c_{k}n_{b}^{-1/10}$. At each perturbation $h_{k}$, we generate $n_{b}$ samples $\bm{\Delta}(h_{k})=\{\Delta_{1}(h_{k}),...,\Delta_{n_{b}}(h_{k})\}$. Repeat the bootstrap procedure in Section 3 independently for $I$ times, and then we obtain $I$ versions of $\widehat{\Delta}^{b}_{n_{b},h_{k}}$, denoted by $\left\{\widehat{\Delta}^{b}_{n_{b},h_{k}}(1),...,\widehat{\Delta}^{b}_{n_{b},h_{k}}(I)\right\}$. Then $\mathbb{E}_{*}\widehat{\Delta}_{n_{b},h_{k}}$ is approximated by
@@ 5.1 Estimating $B$ and $\sigma^{2}(\theta_{0})$
It is worth noting that both of the regression problems (14) and (16) exhibit heteroscedasticity. This arises from the fact that elements of $\mathcal{E}_{e}$ in (15) and $\mathcal{E}_{v}$ in (17) possess varying variances, due to different values of $h_{k}$’s. The primary approach to addressing heteroscedasticity is the weighted least squares method. This involves determining appropriate weights to eliminate or mitigate the heteroscedasticity in the regression models. Because the estimator obtained through the weighted least squares method exhibits the same convergence rate as that derived from the ordinary least squares method (Hayashi, 2011), we omit an in-depth analysis of the weighted least squares method.
@@
However, when estimating $\sigma^{2}(\theta_{0})$, we lack such standard deviations. From $\mathcal{E}_{v}$ in (17), it can be seen that the heteroscedasticity of the regression model (16) arises from the different variances of $\phi(\cdot)$ for different values of $h_{k}$’s. It follows from (13) that $h_{k}^{2}\phi(h_{k})$ ($k=1,...,K$) share the same variance. Therefore, we can handle the heteroscedasticity resulting from differences in perturbations by multiplying both sides of (16) by $h_{k}^{2}$ ($k=1,...,K$) simultaneously. That is, we regress $\widetilde{\bm{Y}}_{v}$ on $\widetilde{\bm{X}}_{v}$, where
@@ 5.2 Estimating $\alpha^{\prime}(\theta_{0})$
Once we have generated $\Delta_{j}(h_{k})$, where $j=1,2,...,n_{b}$ and $k=1,2,...,K$, and estimated the unknown constants, we can transform these samples using (31) and construct the Cor-CFD estimator using (32). We summarize the above procedure as Algorithm 1:
@@
    Usually, we choose $\mathcal{P}_{0}$ as a truncated normal distribution $\mathcal{P}_{0}=\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$, where $0<L<U<\infty$. This setting ensures that $\mathbb{E}[H_{K}]$, $\mathrm{Var}[H_{K}]$, $\mathbb{E}[V_{K}]$, $\mathbb{E}\left[\widetilde{H}_{K}\right]$, $\mathrm{Var}\left[\widetilde{H}_{K}\right]$, $\mathbb{E}\left[\widetilde{V}_{K}\right]$, $\mathbb{E}\left[\widehat{H}_{K}\right]$,$\mathrm{Var}\left[\widehat{H}_{K}\right]$ and $\mathbb{E}\left[\widehat{V}_{K}\right]$ are all bounded. For $c_{1},...,c_{K}\sim\mathcal{P}_{0}$, the condition $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ holds only when $U\leq\sqrt{2}L$, which is a stringent condition. However, we have mentioned that $\max_{k\in\{1,...,K\}}c_{k}^{2}\leq 2\min_{k\in\{1,...,K\}}c_{k}^{2}$ is a sufficient condition of $q\leq K$ which is satisfied more easily for $\mathcal{P}_{0}$ with different parameters.
@@
    In step 1, the pilot samples are generated at the perturbations $h_{k}=c_{k}n_{b}^{-1/10}$, rather than $c_{k}n_{b}^{-1/6}$. This is because (23) and (24) indicate that the optimal perturbation for the pilot samples is of order $n_{b}^{-1/10}$.
@@
    In step 2, we generate the remaining samples at $\widehat{h}_{n}$, instead of $\widehat{h}_{n_{2}}$ defined as in (25), because we will utilize the total $n$ samples to construct the CFD estimator.
Input: The number of total sample pairs $n$, the number of perturbation parameters $K$, the number of pilot sample pairs at each perturbation $n_{b}$, the number of bootstrap $I$ and initial perturbation generator $\mathcal{P}_{0}$.
@@
    Regress $\widetilde{\bm{Y}}_{e}$ on $\widetilde{\bm{X}}_{e}$, where $\widetilde{\bm{Y}}_{e}$ and $\widetilde{\bm{X}}_{e}$ are defined in (43), and the slope is $\widehat{B}$.
    Regress $\widetilde{\bm{Y}}_{v}$ on $\widetilde{\bm{X}}_{v}$, where $\widetilde{\bm{Y}}_{v}$ and $\widetilde{\bm{X}}_{v}$ are defined in (44), and the slope is $\widehat{\bm{\beta}}_{v}$.
@@
Step 2. Construct the Cor-CFD estimator of $\alpha^{\prime}(\theta_{0})$.
@@
    Reuse $\Delta_{j}(h_{k})$, where $j=1,2,...,n_{b}$ and $k=1,2,...,K$. That is, calculate $\Delta_{j}^{cor}(h_{k})=\dfrac{|h_{k}|}{\left|\widehat{h}_{n}\right|}\left(\Delta_{j}(h_{k})-\left(1,h_{k}^{2}\right)\widehat{\bm{\beta}}_{e}\right)+\left(1,{\widehat{h}}_{n}^{2}\right)\widehat{\bm{\beta}}_{e}$ for $j=1,2,...,n_{b}$, and $k=1,2,...,K$.
@@
    Generate $n_{2}=n-Kn_{b}$ sample pairs at $\widehat{h}_{n}$ and calculate $\left\{\Delta_{1}(\widehat{h}_{n}),...,\Delta_{n_{2}}(\widehat{h}_{n})\right\}$.
@@
    The Cor-CFD estimator
    $\widehat{\Delta}_{n,\widehat{h}_{n}}=\frac{1}{n}\left(\sum_{i=1}^{n_{2}}\Delta_{i}\left(\widehat{h}_{n}\right)+\sum_{k=1}^{K}\sum_{j=1}^{n_{b}}\Delta_{j}^{cor}(h_{k})\right).$
@@ 6 Numerical Experiments
In this section, we apply Algorithm 1 to three examples to verify theoretical results, including the constants estimation and the asymptotic variance reduction and possible bias reduction of our proposed Cor-CFD estimator. All numerical results are based on 1000 replications and additional results are provided in Appendix C.
Across all examples, when applying Algorithm 1, we set the initial perturbation generator $\mathcal{P}_{0}=\psi\left(0,1,0.1,\infty\right)$, the number of bootstrap $I=1000$, and the number of regression points $K=10$ or $20$.
@@
Consider estimating the first order derivative of the function $\alpha(\theta)=\kappa\cdot{\rm sin}(\theta)$ at the point $\theta_{0}=0$, where $\kappa$ is a priori unknown parameter.
@@
$\alpha^{\prime}(\theta_{0})=\kappa,\quad B=-\kappa/6,\quad\sigma^{2}(\theta_{0})=1,$
@@
Example 1, from Li and Lam (2020), is used to assess the validity and robustness of the sample-driven method for estimating unknown constants and perturbations. We compare the performance of our proposed method and the estimation-minimization CFD (EM-CFD) method proposed by Li and Lam (2020). In the implementation, we set $\kappa=10$, $K=10$ and utilize 200 pilot sample pairs. The results, presented in Figures 1 and 2, display three subfigures from left to right, representing the estimators of $B$, $\sigma^{2}(\theta_{0})$, and $h$ across 1000 replications. We compare the performance of our proposed estimator and the estimation-minimization CFD (EM-CFD) estimator proposed by Li and Lam (2020). When executing the EM-CFD algorithm, to align with the setting in Li and Lam (2020), we generate $h_{1},\ldots,h_{200}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}\left(0,1\times 200^{-1/5}\right)$.
@@
Figure 1: Comparison of the estimates of the unknown constants and perturbation for case 1 in Example 1.
In the left subfigure of Figure 1, we observe that our method’s robustness in estimating the bias constant is slightly lower than that of the EM-CFD algorithm in case 1, as indicated by the wider box for the Cor-CFD method. This could be due to the larger number of regression points used in the EM-CFD method; the Cor-CFD method employs only $K=10$ regression points, whereas EM-CFD uses 200 sample pairs as regression points. Nonetheless, the two estimators of the bias constant are still comparable overall, likely due to the reduced error at each regression point in our proposed sample-driven method. The middle subfigure of Figure 1 indicates that our method effectively estimates the variance constant, a challenging task for the EM-CFD algorithm as it incorporates the variance of the perturbation generator. The right subfigure provides a comprehensive view of perturbation estimation. Across all 1000 experiments, the estimated perturbations through our method are much closer to the true value.
@@
In Figure 2, we illustrate the impact of different methods on the estimation of $B$, $\sigma^{2}(\theta_{0})$, and $h$ in case 2 of Example 1, which deals with scenarios of heteroskedasticity. Analyzing the pair of subfigures presented on the left, it is apparent that our method captures the variance information effectively and provides a more robust estimator of the bias constant. From the right subfigure, we observe that our method yields a more accurate and reliable estimator of the perturbation.
Furthermore, in Appendix C.1, we verify the sensitivity of different methods w.r.t. $r$, demonstrating the importance of inducing correlated samples in our method.
@@
Figure 2: Comparison of the estimates of the unknown constants and perturbation for case 2 in Example 1.
@@
We consider estimating the first order derivatives of the polynomial function $\alpha(\theta)=1-6\theta+6\theta^{2}-2.5\theta^{3}+0.1\theta^{5}$ at different $\theta_{0}$’s. Assume that the observed variables at $\theta$ obey a normal distribution with mean $\alpha(\theta)$ and variance 1, i.e., $Y(\theta)\sim\mathcal{N}(\alpha(\theta),1)$. From a simple calculation we obtain
$\alpha^{\prime}(\theta_{0})=-6+12\theta_{0}-7.5\theta_{0}^{2}+0.5\theta_{0}^{4},\quad B=-2.5+\theta_{0}^{2},\quad\sigma^{2}(\theta_{0})=1,$
@@
In this example, we attempt to verify the asymptotic variance reduction and possible bias reduction of our proposed Cor-CFD estimator. Table 3 presents a comparison of bias, variance, and MSE between the Cor-CFD estimator (denoted by Bias-Cor, Var-Cor and MSE-Cor, respectively) and the Opt-CFD estimator (denoted by Bias-Opt, Var-Opt and MSE-Opt, respectively) across various sample-pair sizes and $\theta_{0}$’s in Example 2. Additional experiment results are provided in Appendix C.2, comparing the estimated MSEs of different methods. In Table 3, $B$ and $D$ are respectively the true values and Opt-CFD is meant to set $B$ and $\sigma^{2}(\theta_{0})$ as their true values, i.e., $B=-2.5+\theta_{0}^{2}$ and $\sigma^{2}(\theta_{0})=1$. In addition, $\Lambda=\bm{c^{\top}Pc^{4}}\approx-2.2450$. We take the total samples to estimate the unknown constants (i.e., $r=1$) and set $K=10$ for the Cor-CFD estimator.
Table 3: Comparison of the bias, variance and MSE of the Cor-CFD and Opt-CFD estimators in Example 2.
Budget $x$ $B$ $D$ Bias-Cor Bias-Opt Var-Cor Var-Opt MSE-Cor MSE-Opt
$10^{2}$ 0 $-2.5$ 0.1 $-0.3034$ $-0.1851$ 0.0711 0.0655 0.1631 0.0997
1 $-1.5$ 0.1 $-0.2616$ $-0.1581$ 0.0679 0.0489 0.1362 0.0739
2 1.5 0.1 0.0351 0.1501 0.0658 0.0474 0.0670 0.0699
3 6.5 0.1 0.1673 0.2340 0.0750 0.1313 0.0134 0.1029
$10^{3}$ 0 $-2.5$ 0.1 $-0.1256$ $-0.0865$ 0.0111 0.0131 0.0269 0.0205
1 $-1.5$ 0.1 $-0.1169$ $-0.0670$ 0.0099 0.0100 0.0235 0.0145
2 1.5 0.1 0.0345 0.0757 0.0099 0.0098 0.0111 0.0155
3 6.5 0.1 0.0889 0.1150 0.0117 0.0291 0.0196 0.0423
$10^{4}$ 0 $-2.5$ 0.1 $-0.0559$ $-0.0380$ 0.0018 0.0034 0.0049 0.0048
1 $-1.5$ 0.1 $-0.0487$ $-0.0352$ 0.0016 0.0023 0.0040 0.0035
2 1.5 0.1 0.0157 0.0309 0.0017 0.0023 0.0019 0.0033
3 6.5 0.1 0.0414 0.0521 0.0021 0.0061 0.0038 0.0088